# EMOVA[Uncaptioned image]: Empowering Language Models to See, Hear and Speak with Vivid Emotions

### v2

@@ Appendix E More on Vision-language
This section analyzes the pre-training configurations of the vision encoder and the prompt template during Stage 2, as shown in Table. 7. Our final selection is the colored setting. We find that training the ViT model with half of the deeper parameters (Chen et al., 2023d) with a learning rate of 2×$\times 10^{-6}$ (Liu et al., 2024a) yields the best performance. Furthermore, we compare the QA template with the plain template (Liu et al., 2024b) and find that the QA template is superior for pre-training.
@@ Appendix F Limitations
In the current version, EMOVA can only process either visual/speech/text inputs or produce speech/text outputs at the same time. For a communication experience that mirrors human interaction, handling inputs and outputs simultaneously is crucial. Recent works like VITA (Fu et al., 2024b) and LSLM (Ma et al., 2024) have begun to explore duplex modeling. VITA focuses on recognizing speech in noisy environments during the generation process to facilitate timely responses. LSLM, on the other hand, attempts to halt speech production when it detects a command or voice. Recently, a ground-breaking work named Moshi (Défossez et al., 2024) develops a model supporting fully duplex modeling. The adeptness at simultaneously managing the information streams from both the user and the assistant allows Moshi to converse with human beings in real-time scenarios.
However, incorporating emotion into this duplex modeling presents additional challenges. Emotional states can fluctuate throughout a conversation, and how to generate appropriate responses given the user’s previous and current emotional cues has not been thoroughly investigated. We will dive into this topic in the future work.
@@
Although speech units have served as speech representation, they are predominantly adopted in conjunction with text-based assistance (Zhang et al., 2023a). However, the direct generation from unit to unit without text assistance is an area that has not been extensively explored. In Lee et al. (2021), speeches from the source language are directly translated into speech units of the target language for speech-to-speech translation. Similarly, Nguyen et al. (2023) builds a language model directly on speech units, enabling spoken dialogue generation from raw audio. Both works develop models in speech-only data.
@@
Currently, we only utilize a single vision encoder pre-trained via a vision-language manner, while recent works have shown effectiveness by combining vision encoders pre-trained by different manners (e.g., self-supervised pre-training (Liu et al., 2022)) and architectures (e.g., MoE (Zhili et al., 2023; Liu et al., 2024c)). We prioritize visual understanding in this work, while the incorporation of (controllable) visual generation (Chen et al., 2023c; Gao et al., 2023; Li et al., 2023b; Wang et al., 2024; Liu et al., 2023d; Gao et al., 2024) is also appealing to better empower EMOVA in real-life applications (Li et al., 2024c; 2022; Han et al., 2021).
@@ Appendix G Qualitative Results
EMOVA exhibit advanced vision-language abilities in humor understanding (Fig. 11), numerical calculations (Fig. 12), coding (Fig. 13), geometry problem solving (Fig. 15).
@@
EMOVA can engage in omni-modal emotional spoken dialogue (Figs. 6, 8, 14, 17). For demonstration purposes, we present only the speech transcriptions [^2] For speech files, please refer to our project page. of the instruction and EMOVA’s response. The text highlighted in blue indicates the style labels predicted by EMOVA, reflecting the emotion and pitch of the generated speech.
@@
EMOVA can perform structural data understanding in spoken dialogue. The data types include PDF files (Fig. 5), movie posters (Figs. 16 and 18), personal résumés (Fig. 19), charts (Fig. 20), and websites (Fig. 21). Remarkably, EMOVA is not specifically trained on spoken dialogue involving these data types. This suggests that our chain of modality data introduced in Sec. 4.2 effectively activates both vision-language abilities and speech understanding acquired during the text-centric alignment stage.

### v3

@@ Appendix E More on Vision-language Architecture
This section analyzes the pre-training configurations of the vision encoder and the prompt template during Stage 2, as shown in Table. 7. Our final selection is the colored setting. We find that training the ViT model with half of the deeper parameters [9] with a learning rate of 2×$\times 10^{-6}$ [48] yields the best performance. Furthermore, we compare the QA template with the plain template [49] and find that the QA template is superior for pre-training.
@@ Appendix F Limitations
In the current version, EMOVA can only process either visual/speech/text inputs or produce speech/text outputs at the same time. For a communication experience that mirrors human interaction, handling inputs and outputs simultaneously is crucial. Recent works like VITA [19] and LSLM [60] have begun to explore duplex modeling. VITA focuses on recognizing speech in noisy environments during the generation process to facilitate timely responses. LSLM, on the other hand, attempts to halt speech production when it detects a command or voice. Recently, a ground-breaking work named Moshi [13] develops a model supporting fully duplex modeling. The adeptness at simultaneously managing the information streams from both the user and the assistant allows Moshi to converse with human beings in real-time scenarios.
However, incorporating emotions into this duplex modeling presents additional challenges. Emotional states can fluctuate throughout a conversation, and how to generate appropriate responses given the user’s previous and current emotional cues has not been thoroughly investigated. We will dive into this topic in the future work.
@@
Although the speech units have served as the speech representation, they are predominantly adopted in conjunction with text-based assistance [94]. However, the direct generation from unit to unit without text assistance is an area that has not been extensively explored. In [37], speeches from the source language are directly translated into speech units of the target language for speech-to-speech translation. Similarly, [68] builds a language model directly on speech units, enabling spoken dialogue generation from raw audio. Both works develop models in speech-only data.
@@
Currently, we only utilize a single vision encoder pre-trained via a vision-language manner, while recent works have shown effectiveness by combining vision encoders pre-trained by different manners (e.g., self-supervised pre-training [52]) and architectures (e.g., MoE [97, 54]). We prioritize visual understanding in this work, while the incorporation of (controllable) visual generation [8, 22, 42, 83, 53, 23, 24] is also appealing to better empower EMOVA in real-life applications [44, 41, 29, 85]. Digging into its robustness towards noisy vision inputs [28] is also an appealing direction.
Comparison with Emu3 [82].
Both our EMOVA and Emu3 build end-to-end MLLMs with discrete tokenization. Emu3 discretizes visual data, enabling visual understanding and generation, while our EMOVA discretizes speech data with a continuous visual encoder, building an Omni-modal LLM with visual, text, and speech abilities simultaneously.
@@ Appendix G Qualitative Results
Our EMOVA exhibits advanced vision-language abilities in humor understanding (Fig. 10), numerical calculations (Fig. 11), coding (Fig. 12), geometry problem solving (Fig. 16).
@@
EMOVA can engage in omni-modal emotional spoken dialogue (Figs. 5, 13, 14, 18). For demonstration purposes, we present only the speech transcriptions [^2] For speech files, please refer to our project page. of the instruction and EMOVA’s response. The text highlighted in blue indicates the style labels predicted by EMOVA, reflecting the emotion and pitch of generated speeches.
@@
EMOVA can perform structural data understanding even in spoken dialogue. The data types include PDF files (Fig. 15), movie posters (Figs. 17 and 19), personal résumés (Fig. 20), charts (Fig. 21), and websites (Fig. 22). Remarkably, EMOVA is not specifically trained on spoken dialogue data involving these data types. This suggests that our chain of modality data introduced in Sec. 4.2 effectively activates both vision-language abilities and speech understanding acquired during the text-centric alignment stage.
@@
Figure 13: EMOVA engages in spoken dialogue with a cheerful tone.